This is not in a textbook it's in a Math Student workbook made up by my university.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Jennie drives to her son's house in 5 hours when there is heavy traffic. When there is no traffic, the same trip takes 4 hours. She can drive an average of 12 m.p.h. faster when there is no traffic than when thereis heavy traffic.
What is her average speed when there is no traffic?
What is the distance from her house to her son's house?

Answer by ankor@dixie-net.com(12701)

Jennie drives to her son's house in 5 hours when there is heavy traffic. When there is no traffic, the same trip takes 4 hours. She can drive an average of 12 m.p.h. faster when there is no traffic than when there is heavy traffic.
:
What is her average speed when there is no traffic?
:
Let s = Av speed, no traffic
Then
(s-12) = Av speed in traffic
:
The distance for both situations is given as the same:
:
Write a distance equation: Distance = time * speed
:
Dist traffic = Dist no traffic
5(s-12) = 4s
5s - 60 = 4s
5s - 4s = +60
s = 60 mph no traffic
and
60 - 12 = 48 mph in traffic
:
:
What is the distance from her house to her son's house?
5(48) = 240 miles
and
4(60) = 240 miles